Curriculum Vitae for

Bozhidar Zakhariev Iliev^{∗}^{†}

NAME: Bozhidar Zakhariev ILIEV (this is the English transliteration; the French one is Bojidar Zahariev ILIEV)

NATIONALITY and CITIZENSHIP: Bulgarian

DATE OF BIRTH: 18 April, 1958

PLACE OF BIRTH: Sofia, Bulgaria

LANGUAGES: Bulgarian (native speaker), Russian (fluent), English (good), German (bad)

OFFICE ADDRESS: Laboratory of Mathematical Modeling in Physics, Institute for Nuclear Research and Nuclear Energy, Bulgarian Academy of Sciences, Boul. Tzarigradsko chaussée 72, 1784 Sofia, Bulgaria; Tel: +359-2-7144-612, Fax: +359-2-9753619

MARITAL STATUS: divorced, 1 daughter (Victoria, b. 18 February 1991)

EDUCATION:

• HIGH SCHOOL: 1976 (Jun 27), The National Mathematical School in Sofia.

• UNIVERSITY: 1981 (Mar 23), The Physical faculty of the Sofia State University, qualification physicist with specialization atomic physics (degree: M. Sc. (BSc, Dipl. phys.) in physics).

• DEGREES: 1995 (Mar 7), Ph.D. in mathematics, Institute of Mathematics of the Bulgarian Academy of Sciences. Advisers of the thesis, “Transports along paths in fibre bundles”, were prof. Stancho Dimiev and prof. Vladimir Aleksandrov from the division of Complex Analysis.

ACADEMIC RANK:

• 1981 (Jul 1) – 1986 (Dec 31): Physicist, M. Sc.

• 1988 (Jan 1) – 1996 (Mar 7): Physicist, M. Sc.

• 1996 (Mar 7) – 1999 (Apr 29): Assistant Professor (Research fellow) II degree, Ph.D. (in math.)

• 1999 (Apr 29) – 2002 (Nov 19): Assistant Professor (Research fellow ) I degree, Ph.D. (in math.)

• Since 2002 (Nov 19): Associate Professor II degree, Ph.D. (in math.)

∗E-mail address: bozho@inrne.bas.bg †URL: http://theo.inrne.bas.bg/∼bozho/

PROFESSIONAL DISTINCTIONS:

• Gold medal from the National Mathematical School in Sofia, June, 1976.

• Included in “Who’s Who in the World”, 2000 Millennium Edition, A Who’s Who in America Publication, Marquis Who’s Who, New Providence, January 2000.

• Included in “Outstanding people of the 20th Century”, Second edition, International Biographical Centre, Cambridge, England, 2000.

• Included in “2000 outstanding intellectuals of the 20th century”, International Biographical Centre, Cambridge, England, 2000; Second edition, 2003.

• Included in “2000 outstanding scientists of the 21st century”, 1st Edition, International Biographical Centre, Cambridge, England, 2001.

• Included in “Whos who in the 21st century”, International Biographical Centre, Cambridge, England, 2001.

• Included in “Who’s Who in the World, 2001”, 18th Edition, A Who’s Who in America Publication, Marquis Who’s Who, New Providence, 2001.

• Included in “Dictionary of international biography”, 29th Edition, International Biographical Centre, Cambridge, England, 2001.

• Included in “Who’s Who in Science and Engineering, 2002∼2003”, 6th Edition, A Who’s Who in America Publication, Marquis Who’s Who, New Providence, 2003.

• Included in ”Who’s Who”, American Biographical Institute (Raleigh, North Carolina 27622, USA), 2004.

• Included in ”The Contemporary Who’s Who of Professionals”, American Biographical Institute (Raleigh, North Carolina 27622, USA), 2005.

ADDITIONAL ASSIGNMENTS:

• Invited Participant at the International Conferences: “Gravitational Energy and Gravitational Waves”, Joint Institute for Nuclear Research, Dubna USSR (Russia): 1989, 1990, 1991, 1992, 1993, 1994, 1995, 1996.

• Visitor: College on Differential Geometry (from 30.10.1989 to 01.12.1989), International Centre for Theoretical Physics in Trieste, Italy.

• Invited Participant at the International Workshop “Advanced Geometric Methods in Physics”, Florence, Italy, 14 – 18 April, 2005.

• Invited Participant at the International Workshop on “Complex Structures and vector Fields”:

- 6th, September 3 – 7, 2002, Golden Sends resort (near Varna), Bulgaria.

- 7th, August 28 – September 3, 2004, Plovdiv, Bulgaria.

- 8th, August 21 – 26, 2006, Sofia, Bulgaria

- 9th, August 25 – 29, 2008, Sofia, Bulgaria.

FIELDS OF PROFESSIONAL ACTIVITY:

• Differential Geometry

• General Relativity and its generalizations

• Quantum Mechanics

• Quantum Field Theory

• (Particle Physics)

MAIN TOPICS OF PROFESSIONAL ACTIVITY:

• Quantum field theory

• Quantum mechanics on fibre bundles

• The equivalence principle (in gravity physics)

• Normal frames and coordinates

• Transports along paths in fibre bundles and their physical

• Riemannian metrics

• Connection theory

• Deviation equations in spaces with affine connection applications

JOB EXPERIENCE:

• Physicist, assistant prof. and associated prof. in the Institute for Nuclear Research and Nuclear Energy of the Bulgarian Academy of Sciences. From 01.07.1981 to 31.12.1986 and since 01.01.1988.

• Assistant professor in the group of Gravity Physics and Relativity Theory (of prof. N. A. Chernikov) in the Bogoliubov Laboratory of Theoretical Physics and in the Laboratory of Computing Technique and Automation of the Joint Institute for Nuclear Research in Dubna (USSR, now Russia):

– from 30.11.1981 to 17.12.1981; from 25.10.1983 to 26.01.1984;

– from 03.12.1986 to 03.05.1987; from 27.11.1988 to 09.12.1988;

– from 18.07.1989 to 18.10.1989; from 09.03.1990 to 11.04.1990;

– from 14.05.1990 to 23.05.1990; from 01.08.1990 to 15.12.1990;

– from 10.11.1992 to 14.01.1993; from 05.06.1993 to 14.08.1993;

– from 29.12.1993 to 14.02.1994; from 18.04.1994 to 31.05.1994;

– from 30.01.1995 to 28.03.1995; from 26.11.1996 to 27.02.1997.

• Scientific visits (field trips) in:

– International Centre for Theoretical Physics in Trieste, Italy (College on Differential Geometry; from 30.10.1989 to 01.12.1989).

– Novosibirsk State University, Novosibirsk, Russia (former USSR)(from 19.02.1990 to 02.03.1990).

– Yaroslav State University, Yaroslav, Russia (from 13.06.1993 to 19.06.1993).

– Tartu University, Tartu, Estonia (from 26.06.2004 to 16.07.2004 and from 27.05.2005 to 07.06.2005).

SCIENTIFIC PUBLICATIONS

• More than 80 titles, including:

- 2 monographs

- more than 30 papers in specialized journals

- 25 JINR Communications

- 7 chapters in books

- more than 60 electronic preprints

MEMBER OF THE EDITORIAL BOARD OF

• International Journal of Geometric Methods in Modern Physics

MEMBER OF THE ORGANIZING COMMETEE OF

• 8th International Workshop on “Complex Structures and vector Fields”, August 21 – 26, 2006, Sofia, Bulgaria.

• 9th International Workshop on “Complex Structures, Integrability and vector Fields”, August 25 - 29, 2008, Sofia, Bulgaria.

OTHER CREDENTIALS:

• 1986: Master SCUBA diver

• 1988: Skipper of seagoing vessel up to 20 GT

• 1989: Driving licence categories B & M

• 1996: Diving instructor

• 1998: SCUBA diving instructor

Date: June 2, 2009 Signature:

/B. Z. Iliev/